School starts at 8:00 AM. A bell rings every 60 minutes, and another bell rings every 48 minutes. If both bells ring at the start of the day, how many minutes will it be before they ring at the same time again? Show all work.
Answer by ikleyn(52781)

.

You should find the Least Common Multiple of the numbers 60 and 48.


    60 = 12*5;  48 = 12*4;  therefore, LCM(60,48) = 12*4*5 = 240.


ANSWER.  The two bells ring simultaneously every 240 minutes, i.e. every 4 hours.

Solved.   //   And all relevant key terms/words were pronounced.